# Bright and dark breathers on an elliptic wave in the defocusing mKdV equation

### v1

@@ Abstract.
Breathers on an elliptic wave background consist of nonlinear superpositions of a soliton and a periodic wave, both traveling with different wave speeds and interacting periodically in the space-time. For the defocusing modified Korteweg–de Vries (mKdV) equation, the construction of general breathers has been an open problem since the elliptic wave is related to the elliptic degeneration of the hyperelliptic solutions of genus two. We have found the new representation of eigenfunctions of the Lax operator associated with the elliptic wave, which enables us to solve this open problem and to construct two families of breathers with bright (elevation) and dark (depression) profiles.
@@ 1. Introduction
We consider the integrable model given by the defocusing modified Korteweg–de Vries (mKdV) equation written in the form:
@@
where $(x,t)\in\mathbb{R}\times\mathbb{R}$ and $u=u(x,t)\in\mathbb{R}$. As is well-known (see reviews in [18, 19]), there exists a three-parameter family of the traveling periodic wave solutions of the defocusing mKdV equation (1.1), which we call the elliptic wave. In the symmetric case, it reduces to the snoidal profile, which is related to the genus-one elliptic solutions of integrable equations of the AKNS hierarchy [4]. In the non-symmetric case, it is related to the elliptic degeneration of the genus-two hyperelliptic solutions. Although Riemann theta functions degenerate into Jacobi theta functions in this limit [20], the general theory is too abstract for practical applications of the mKdV equation.
The mKdV equation is one of the fundamental models of the dispersive hydrodynamics, where the elliptic waves and breathers on their backgrounds capture dynamics of dispersive shock waves and soliton gases [1]. Breathers are nonlinear superpositions of a soliton and a periodic wave, both traveling with different wave speeds and interacting periodically in the space-time. In the symmetric case, breathers with the dark (depression) profiles were constructed in [37] by using the explicit expression for eigenfunctions of the Lax operator associated with the genus-one elliptic wave from [40]. In the non-symmetric case, breathers with the kink profiles were constructed in [3] for the only case when the spectral parameter is at the origin (which is the center of symmetry of the Lax spectrum). The main purpose of this work is to construct breathers with both bright (elevation) and dark (depression) profiles in the general case of the genus-two elliptic wave and for the arbitrary non-zero values of the spectral parameter.
@@ 1.1. Main results
The general periodic solution of the system (1.4), (1.5), and (1.6) has different analytic forms for $b=0$ (symmetric case) and $b\neq 0$ (non-symmetric case). If $b\neq 0$ (without loss of generality, we consider $b>0$), the general periodic solution is best parameterized by the real parameters $(\zeta_{1},\zeta_{2},\zeta_{3})$ satisfying $0<\zeta_{3}<\zeta_{2}<\zeta_{1}$ in the form:
@@
with two symmetric bandgaps $(-\zeta_{1},-\zeta_{2})\cup(\zeta_{2},\zeta_{1})$ and the central bandgap $(-\zeta_{3},\zeta_{3})$. If $0<\zeta_{3}<\zeta_{2}<\zeta_{1}$, then the wave profile $\phi$ is periodic on $\mathbb{R}$ with the period $2\nu^{-1}K(k)$ and on $i\mathbb{R}$ with the period $2\nu^{-1}K^{\prime}(k)$. In what follows, we drop the dependence of the elliptic functions and integrals on $k\in(0,1)$ if it does not cause a confusion.
@@
Figure 1 shows the pre-image of the mapping $0,K]\times[0,iK^{\prime}]\ni\gamma\to\zeta\in[0,\infty)$, based on Lemma 1 below. With this parameterization, the following theorem specifies the explicit representation of the eigenfunctions.
@@
$\left(\begin{matrix}p_{1}(x)\\ q_{1}(x)\end{matrix}\right)=\left(\begin{matrix}-i\zeta\frac{H(\nu x+\gamma+\alpha)}{\Theta(\nu x+\alpha)}+\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)}\frac{H(\nu x+\gamma-\alpha)}{\Theta(\nu x-\alpha)}\\ -i\zeta\frac{H(\nu x+\gamma+\alpha)}{\Theta(\nu x+\alpha)}-\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)}\frac{H(\nu x+\gamma-\alpha)}{\Theta(\nu x-\alpha)}\end{matrix}\right)e^{-\nu xZ(\gamma)},$ (1.16)
@@
1.2. Contruction of breathers
We use the Darboux transformation
@@ 1.2. Contruction of breathers
We show in Section 4 below that if $\zeta\in(0,\zeta_{3})\cup(\zeta_{2},\zeta_{1})$, then we can choose scaling factors in $(p_{1},q_{1})$ and $(p_{2},q_{2})$ such that
@@
This implies that the solution $u=u(x,t)$ in (1.26) is real if $c_{1}c_{2}\in i\mathbb{R}$, $c_{1}^{2},c_{2}^{2}\in\mathbb{R}$. The exponential factors in (1.16) and (1.17) are defined differently between $(0,\zeta_{3})$ and $(\zeta_{2},\zeta_{1})$ by the parameter $\kappa>0$:
@@
It follows from (1.26) that only the factor $c_{2}/c_{1}$ affects the representation of the solutions. To reduce the number of parameters, we introduce the phase translation of the soliton relative to the elliptic wave and define
@@
where the negative sign of $c_{2}$ was chosen to obtain bounded breather solutions. It follows from (1.15) with (1.16) and (1.17) that the breather propagates along the coordinate $\eta=x+c_{s}t+x_{0}$ with the (left-propagating) wave speed $c_{s}$ given by
@@
In comparison with the three-parameter family of the elliptic wave with the profile $\phi$ in (1.8), the breather solution obtained from (1.25), (1.26), and (1.30) has two additional parameters $\zeta\in(0,\zeta_{3})\cup(\zeta_{2},\zeta_{1})$ (which determines uniquely $\gamma\in[0,K]\times[0,iK^{\prime}]$, see Figure 1) and $x_{0}\in\mathbb{R}$. Since $\kappa>0$ in (1.29), it follows from (1.16) and (1.17) that the elliptic wave is shifted across the breather with the phase shift in $x$ given by $2\nu^{-1}(\gamma-\alpha)$.
To illustrate the breather solutions, let us first show the periodic profiles of $\phi$ in Figure 2 for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$ (left) and $(\zeta_{1},\zeta_{2},\zeta_{3})=(1.25,1,0.5)$ (right). In both cases, the difference between (1.8) and (1.12) is found within the machine precision error. The left panel corresponds to $\beta\in(0,K)$, for which $\phi$ has zeros on $\mathbb{R}$, and the right panel corresponds to $\beta\in i(0,K^{\prime})$, for which $\phi$ has no zeros on $\mathbb{R}$.
@@
The breather solution $u=u(x,t)$ in (1.26) for the choice of $\zeta$ in $(0,\zeta_{3})$ is shown for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$ in Figure 3 and for $(\zeta_{1},\zeta_{2},\zeta_{3})=(1.25,1,0.5)$ in Figure 4. In both cases, the left panels show the snapshot versus $x$ at $t=0$ and the right panels show the solution surface versus $(\xi,t)$, where $\xi=x+ct$. In both cases, we have set $x_{0}=0$ in (1.30) and chosen $\zeta$ from (1.14) with $\gamma=2\alpha+0.1(K-2\alpha)+iK^{\prime}$. For the choice of $\zeta$ in $(0,\zeta_{3})$, breathers represent the elevation (bright) profile propagating on the traveling elliptic wave to the right in the reference frame $\xi=x+ct$ moving with the wave speed $c$ since $c_{s}<c$ in (1.31).
In the limit $\gamma\to 2\alpha+iK^{\prime}$ (when $\zeta\to 0$), the breather profile becomes wide and the solution degenerates to the kink breather constructed in [3, Theorem 2]. There are two kinks by the symmetry: from negative to positive values of the periodic wave and from positive to negative values of the periodic wave (the latter is shown in Figure 4 in [3]). The kinks also moves to the right in the reference frame $\xi=x+ct$ moving with the wave speed $c$.
Figure 3. The snapshow versus $x$ for $t=0$ (left) and the solution surface versus $(\xi,t)$ (right) for the bright breather with $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$.
@@
The breather solution $u=u(x,t)$ in (1.26) for the choice of $\zeta$ in $(\zeta_{2},\zeta_{1})$ is shown for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$ in Figure 5 and for $(\zeta_{1},\zeta_{2},\zeta_{3})=(1.25,1,0.5)$ in Figure 6. The spectral parameter $\zeta$ is defined by (1.14) with $\gamma=0.5K$. For the choice of $\zeta$ in $(\zeta_{2},\zeta_{1})$, breathers represent the depression (dark) profile propagating on the traveling periodic background to the left in the reference frame $\xi=x+ct$ moving with the wave speed $c$ since $c_{s}>c$ in (1.31).
In the limit $\zeta_{3}\to 0$, the elliptic wave is reduced to the snoidal profile by the Landen transformation [3, Example 3.2]. Dark breathers on the snoidal profile were constructed explicitly in [37, Theorem 1]. The dark breathers also move to the left in the reference frame $\xi=x+ct$ moving with the wave speed $c$. see Figure 1 in [37].
Figure 5. The snapshow versus $x$ for $t=0$ (left) and the solution surface versus $(\xi,t)$ (right) for the dark breather with $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$.
@@ 1.3. Background of the problem
Inspired by the recent studies of the soliton gas in [13, 23, 24], the breather solutions on the elliptic wave background appear to be the central objects for interactions of solitons in the limit of infinitely many solitons. Consequently, the same solutions of the KdV equation were comprehensively studied in two independent works [6, 26]. It is shown in [26] that the bright profiles are related to the spectral parameter chosen below the Lax spectrum and the dark profiles are related to the choice in a spectral gap of a finite length.
@@
For the defocusing cubic NLS equation, dark profiles of breathers were constructed for the snoidal elliptic wave in [39, 40], see also [36]. It was found recently in [27] that both bright and dark profiles of breathers exist for the elliptic wave with a non-trivial phase, which is still expressed by the genus-one elliptic functions. The elliptic wave background is modulationally unstable for the focusing cubic NLS equation and formations of rogue waves and time-decaying and space-periodic breathers complicate dynamics of perturbations [7, 12, 15].
@@
For the focusing mKdV equation, stability of the elliptic profiles and breathers on their background were only developed in the symmetric case [33, 34]. The main obstacle has been again the explicit characterization of the eigenfunctions of the Lax operators. Only recently in [35], the eigenfunctions are constructed by degeneration of the Riemann theta functions of genus two, very similar to the general elliptic theory in [20]. One of the possible application of our work is to construct the explicit eigenfunctions of the Lax operators for the focusing mKdV equation in terms of the elliptic functions.
Finally, we mention another form used to characterize eigenfunctions of the Lax operators related to the elliptic degeneration of the Riemann theta functions of genus two [30, 31]. The elliptic functions are characterized as quotient of products of Jacobi theta functions centered at the poles and zeros of the elliptic function. Although this form can also be obtained for the eigenfunctions of the Lax system (1.2), we show in Section 5 below that the poles and zeros of the elliptic eigenfunctions are defined by using the branch point singularities, and therefore, the factorization formulas are not as explicit as the representation of eigenfunctions in terms of elliptic functions obtained in Theorem 1.

### v2

@@ Abstract.
Breathers on an elliptic wave background consist of nonlinear superpositions of a soliton and a periodic wave, both traveling with different wave speeds and interacting periodically in the space-time. For the defocusing modified Korteweg–de Vries equation, the construction of general breathers has been an open problem since the elliptic wave is related to the elliptic degeneration of the hyperelliptic solutions of genus two. We have found a new representation of eigenfunctions of the Lax operator associated with the elliptic wave, which enables us to solve this open problem and to construct two families of breathers with bright (elevation) and dark (depression) profiles.
@@ 1. Introduction
We consider the integrable model given by the defocusing mKdV (modified Korteweg–de Vries) equation written in the form:
@@
where $(x,t)\in\mathbb{R}\times\mathbb{R}$ and $u=u(x,t)\in\mathbb{R}$. As is well-known (see reviews in [19, 20]), there exists a three-parameter family of traveling periodic wave solutions of the defocusing mKdV equation (1.1), which we call the elliptic wave. In the symmetric case, it reduces to the snoidal profile, which is related to the genus-one elliptic solutions of integrable equations of the AKNS hierarchy [4]. In the non-symmetric case, it is related to the elliptic degeneration of the genus-two hyperelliptic solutions. Although Riemann theta functions degenerate into Jacobi theta functions in this limit [21], the general theory is too abstract for practical applications of the mKdV equation.
The mKdV equation is one of the fundamental models of the dispersive hydrodynamics, which brings together the balance between the cubic dispersion and the cubic nonlinearity, with many applications in the dynamics of internal waves in seas and oceans [27, 43]. In the dispersive hydrodynamics, the elliptic waves and breathers on their backgrounds capture dynamics of dispersive shock waves and soliton gases [1]. Breathers are nonlinear superpositions of a soliton and a periodic wave, both traveling with different wave speeds and interacting periodically in the space-time. In the symmetric case, breathers with the dark (depression) profiles were constructed in [41] by using the explicit expression for eigenfunctions of the Lax operator associated with the genus-one elliptic functions from [45]. In the non-symmetric case, breathers with the kink profiles were constructed in [3] for the only case when the spectral parameter is at the origin (which is the center of symmetry of the Lax spectrum). The main purpose of this work is to construct breathers with both bright (elevation) and dark (depression) profiles in the general case of the genus-two elliptic functions and for the arbitrary non-zero values of the spectral parameter.
@@ 1.1. Main results
The general periodic solution of the system (1.4), (1.5), and (1.6) has different analytic forms for $b=0$ and $b\neq 0$. For $b=0$ (symmetric case), the level curves of (1.6) on the phase plane $(\phi,\phi^{\prime})$ are symmetric with respect to reflection $\phi\to-\phi$. For $b\neq 0$ (non-symmetric case), the symmetry of reflection $\phi\to-\phi$ is lost on the phase plane $(\phi,\phi^{\prime})$. If $b\neq 0$ (without loss of generality, we consider $b>0$), parameters ($b,c,d$) can be parameterized by the real parameters $(\zeta_{1},\zeta_{2},\zeta_{3})$ satisfying $0<\zeta_{3}<\zeta_{2}<\zeta_{1}$:
@@
with two symmetric bandgaps $(-\zeta_{1},-\zeta_{2})\cup(\zeta_{2},\zeta_{1})$ and the central bandgap $(-\zeta_{3},\zeta_{3})$. If $0<\zeta_{3}<\zeta_{2}<\zeta_{1}$, then the wave profile $\phi$ is periodic on $\mathbb{R}$ with the period $2\nu^{-1}K(k)$ and on $i\mathbb{R}$ with the period $2\nu^{-1}K^{\prime}(k)$, where $K(k)$ is a complete elliptic integral of the first kind and $K^{\prime}(k)=K(k^{\prime})$ with $k^{\prime}=\sqrt{1-k^{2}}$. The spectral stability of the elliptic wave with the profile (1.8) is shown in [15] based on the Lax spectrum (1.10) and the squared eigenfunction relation between eigenfunctions of the Lax system (1.2) and the linearized mKdV equation.
In what follows, we drop the dependence of the elliptic functions and integrals on $k\in(0,1)$ if it does not cause a confusion. We also use standard notations for elliptic functions, which are reviewed in (1.33) and (1.34) below.
@@
This dispersion relation is obtained in Lemma 1 below. Figure 1 shows the pre-image of the mapping $0,K]\times[0,iK^{\prime}]\ni\gamma\to\zeta\in[0,\infty)$, which is justified in Lemma 3 below. With this parameterization, the following theorem specifies the explicit representation of the eigenfunctions. The proof is based on Lemmas 2 and 6 below.
@@
$\left(\begin{matrix}p_{1}(x)\\ q_{1}(x)\end{matrix}\right)=\left(\begin{matrix}-i\zeta\frac{H(\nu x+\gamma+\alpha)}{\Theta(\nu x+\alpha)}+\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)}\frac{H(\nu x+\gamma-\alpha)}{\Theta(\nu x-\alpha)}\\ -i\zeta\frac{H(\nu x+\gamma+\alpha)}{\Theta(\nu x+\alpha)}-\zeta_{1}\frac{\Theta(0)\Theta(2\alpha+\gamma)}{\Theta(\gamma)\Theta(2\alpha)}\frac{H(\nu x+\gamma-\alpha)}{\Theta(\nu x-\alpha)}\end{matrix}\right)e^{-\nu xZ(\gamma)}$ (1.16)
and
@@
1.2. Construction of breathers
Darboux transformation for the mKdV equation are well-known in the literature, e.g., see [28, 40] or [8, Appendix]. The one-fold Darboux transformation is written in the form:
@@ 1.2. Construction of breathers
The main application of Theorem 1 is an algorithm on how to choose $\zeta\in\mathbb{R}$ and $c_{1},c_{2}\in\mathbb{C}$ in order to obtain bounded real-valued solutions $u=u(x,t)$ of the mKdV equation (1.1) from the transformation formula (1.26), which describe breathers on the elliptic wave background. We show in Section 4 below that if $\zeta\in(0,\zeta_{3})\cup(\zeta_{2},\zeta_{1})$, then we can choose scaling factors in $(p_{1},q_{1})$ and $(p_{2},q_{2})$ such that
@@
This implies that the solution $u=u(x,t)$ in (1.26) is real-valued if $c_{1}c_{2}\in i\mathbb{R}$, $c_{1}^{2},c_{2}^{2}\in\mathbb{R}$. The exponential factors in (1.16) and (1.17) are defined differently between $(0,\zeta_{3})$ and $(\zeta_{2},\zeta_{1})$ by the parameter $\kappa>0$:
@@
It follows from (1.26) that only the factor $c_{2}/c_{1}$ affects the representation of the solutions. To reduce the number of parameters, we introduce the phase translation of the breather relative to the elliptic wave and define
@@
where the negative sign of $c_{2}$ was chosen to obtain bounded solutions. The choice of the sign of $c_{2}$ is the only result based on the numerical verifications. All other results have been justified by using elliptic functions.
It follows from (1.15) with (1.16) and (1.17) that the breather propagates along the coordinate $\eta=x+c_{s}t+x_{0}$ with the (left-propagating) wave speed $c_{s}$ given by
@@
In comparison with the three-parameter family of the elliptic wave with the profile $\phi$ in (1.8), the solution obtained from (1.25), (1.26), and (1.30) has two additional parameters $\zeta\in(0,\zeta_{3})\cup(\zeta_{2},\zeta_{1})$ (which determines uniquely $\gamma\in[0,K]\times[0,iK^{\prime}]$, see Figure 1) and $x_{0}\in\mathbb{R}$. Since $\kappa>0$ in (1.29), it follows from (1.16) and (1.17) that the elliptic wave is shifted across the breather with the phase shift in $x$ given by $2\nu^{-1}(\gamma-\alpha)$.
To illustrate the breather solutions, let us first show the periodic profiles of $\phi$ in Figure 2 for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$ (left) and $(\zeta_{1},\zeta_{2},\zeta_{3})=(1.25,1,0.5)$ (right). Either solution form (1.8) or (1.12) can be used for plotting within the machine precision error since the two different representations (1.8) and (1.12) are mathematically equivalent. The left panel corresponds to $\beta\in(0,K)$, for which $\phi$ has zeros on $\mathbb{R}$, and the right panel corresponds to $\beta\in i(0,K^{\prime})$, for which $\phi$ has no zeros on $\mathbb{R}$.
@@
The solution $u=u(x,t)$ in (1.26) for the choice of $\zeta$ in $(0,\zeta_{3})$ is shown for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$ in Figure 3 and for $(\zeta_{1},\zeta_{2},\zeta_{3})=(1.25,1,0.5)$ in Figure 4. In both cases, the left panels show the snapshot versus $x$ at $t=0$ and the right panels show the solution surface versus $(\xi,t)$, where $\xi=x+ct$. In both cases, we have set $x_{0}=0$ in (1.30) and chosen $\zeta$ from (1.14) with $\gamma=2\alpha+0.1(K-2\alpha)+iK^{\prime}$. It is obvious from Figures 3 and 4 that the solution $u=u(x,t)$ represents a breather propagating on the elliptic wave background. For the choice of $\zeta$ in $(0,\zeta_{3})$, the breather has the elevation (bright) profile and it propagates to the right relative to the traveling wave, which is stationary in the reference frame $\xi=x+ct$. This corresponds to $c_{s}<c$ in (1.31).
In the limit $\gamma\to 2\alpha+iK^{\prime}$ (when $\zeta\to 0$), the breather profile becomes wide and the solution degenerates to the kink breather constructed in [3, Theorem 2]. There are two kinks by the symmetry: from negative to positive values of the periodic wave and from positive to negative values of the periodic wave. The kinks also moves to the right relative to the traveling wave, see Figure 4 in [3].
Figure 3. The snapshot versus $x$ for $t=0$ (left) and the solution surface versus $(\xi,t)$ (right) for the bright breather with $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$.
@@
The solution $u=u(x,t)$ in (1.26) for the choice of $\zeta$ in $(\zeta_{2},\zeta_{1})$ is shown for $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$ in Figure 5 and for $(\zeta_{1},\zeta_{2},\zeta_{3})=(1.25,1,0.5)$ in Figure 6. The spectral parameter $\zeta$ is defined by (1.14) with $\gamma=0.5K$. It is again obvious from Figures 5 and 6 that the solution $u=u(x,t)$ represents a breather propagating on the elliptic wave background. For the choice of $\zeta$ in $(\zeta_{2},\zeta_{1})$, the breather has the depression (dark) profile and it propagates to the left relative to the traveling wave, which is stationary in the reference frame $\xi=x+ct$. This coresponds to $c_{s}>c$ in (1.31).
In the limit $\zeta_{3}\to 0$, the elliptic wave is reduced to the snoidal profile by the Landen transformation [3, Example 3.2]. Dark breathers on the snoidal profile were constructed explicitly in [41, Theorem 1]. The dark breathers also move to the left relative to the traveling wave, see Figure 1 in [41].
Figure 5. The snapshot versus $x$ for $t=0$ (left) and the solution surface versus $(\xi,t)$ (right) for the dark breather with $(\zeta_{1},\zeta_{2},\zeta_{3})=(2,1,0.5)$.
@@ 1.3. Background of the problem
Inspired by the recent studies of the soliton gas in [13, 24, 25], the breather solutions on the elliptic wave background appear to be the central objects for interactions of solitons in the limit of infinitely many solitons. Consequently, the same solutions of the KdV equation were comprehensively studied in two independent works [6, 29]. It is shown in [29] that breathers with the bright profiles are related to the spectral parameter chosen below the Lax spectrum and breathers with the dark profiles are related to the choice of the spectral parameter in a spectral gap of a finite length.
@@
For the defocusing cubic NLS equation, dark profiles of breathers were constructed for the snoidal elliptic wave in [44, 45], see also [39]. It was found recently in [30] that both bright and dark profiles of breathers exist for the elliptic wave with a non-trivial phase, which is still expressed by the genus-one elliptic functions. The elliptic wave is modulationally unstable for the focusing cubic NLS equation and formations of rogue waves and time-decaying and space-periodic breathers complicate dynamics of perturbations [7, 12, 16].
@@
For the focusing mKdV equation, stability of the elliptic waves and breathers on their background were only developed in the symmetric case [36, 37]. The main obstacle has been again the explicit characterization of the eigenfunctions of the Lax operators. Only recently in [38], the eigenfunctions are constructed by degeneration of the Riemann theta functions of genus two, very similar to the general elliptic theory in [21]. One of the possible application of our work is to construct the explicit eigenfunctions of the Lax operators for the focusing mKdV equation in terms of the elliptic functions. We anticipate that the eigenfunctions can be obtained in a similar form to the one in Theorem 1 but the challenges arise in analysis of the pre-image of the mapping $0,K]\times[0,iK^{\prime}]\ni\gamma\to\zeta\in\sigma_{L}$ since the Lax spectrum $\sigma_{L}$ is no longer a subset of $\mathbb{R}$.
Finally, we mention another form used to characterize eigenfunctions of the Lax operators related to the elliptic degeneration of the Riemann theta functions of genus two [33, 34]. The elliptic functions can be written in the factorized form given by as quotient of products of Jacobi theta functions, similarly to the forms (1.12) and (1.13). The factorized form for the eigenfunctions of the Lax system (1.2) are expected to be useful in the construction of breathers on the general elliptic wave by using the transformation formula (1.25), as it was the case for the snoidal wave in [41]. However, we show in Section 5 below that the poles and zeros of the elliptic eigenfunctions in the factorized form depend on the spectral parameter via the branch point singularities, and therefore, the factorized form is not as explicit as the representation of eigenfunctions obtained in Theorem 1.